# Entropic Risk Optimization in Discounted MDPs: Sample Complexity Bounds with a Generative Model

### v2

@@ Appendix D Deterioration of Greedy Policy
Appendix E Concentration Results
In this section we collect some lemmas on concentration. they directly complement lemma 3.
Let $N$ denote the number of calls to the generative model on each state-action pair such that the total number of calls is $SAN$. Let $\widehat{P}(s^{\prime}|s,a)$ denote the plug-in estimator obtained from $N$ samples of $s^{\prime}\sim P(\cdot|s,a)$, that is $\widehat{P}(s^{\prime}|s,a)=\frac{\sum_{n=1}^{N}1_{\{s_{n}=s^{\prime}\}}}{N}$ where $s_{n}$ is the outcome of the random variable $X_{n}$ taking values in $\mathcal{S}$ according to $P(\cdot|s,a)$.
Lemma 7.
Fix $\varepsilon>0$. If every state-action pair has been tried at least $N=8\frac{S+\log(\frac{SA}{\delta})}{\tau^{2}}$ times, then it holds that $\max_{(s,a)}\|P(\cdot|s,a)-\widehat{P}(\cdot|s,a)\|_{1}\leq\varepsilon$ with probability at least $1-\delta$.
Proof.
Using the Weissman inequality [57], any confidence interval for a state-action pair that have been tried $m$ times have size $2\sqrt{2[\log(2^{S}-2)-\log(\delta_{P})]/N}$. Setting this size to be smaller than $\tau$ and solving for $N$, we find that
$N\geq\frac{8}{\varepsilon^{2}}\Big(\log(2^{S}-2)-\log(\delta_{P})\Big)\,.$ (60)
Noting that $8\big(S+\log(1/\delta_{P})\big)/\varepsilon^{2}\geq 8[\log(2^{S}-2)-\log(\delta_{P})]/\varepsilon^{2}$ and substituting $\delta_{P}=\frac{\delta}{SA}$, the result follows by the union bound since each of the $SA$ confidence balls contain $P(\cdot|s,a)$ with probability at least $1-\delta$. ∎
Theorem 8.
Let $\pi$ be any fixed policy. For $\beta>0$ we have that if $N>\frac{(1-e^{-\beta\frac{1}{1-\gamma}})^{2}}{2\varepsilon^{2}}\log(2SA/\delta)$ then it holds that with probability at least $1-\delta$ that
$\max_{s,a}\big|\sum_{s^{\prime}}[P(s^{\prime}|s,a)-{\widehat{P}(s^{\prime}|s,a)}]e^{-\beta V^{\pi}(s^{\prime})}\big|<\varepsilon$ (61)
and if $\beta<0$ it holds that if $N>\frac{(1-e^{-\beta\frac{1}{1-\gamma}})^{2}}{2\varepsilon^{2}}\log(2SA/\delta)$, then it holds that with probability at least $1-\delta$ that
$\max_{s,a}\big|\sum_{s^{\prime}}[P(s^{\prime}|s,a)-{\widehat{P}(s^{\prime}|s,a)}]e^{|\beta|(V^{\pi}(s^{\prime})-\frac{1}{1-\gamma})}\big|<\varepsilon\,.$ (62)
Proof.
We only prove the first claim: $\beta>0$ as the other case is completely similar. We note that for the random variable $\sum_{s^{\prime}}1_{\{X_{n}=s^{\prime}\}}(s^{\prime}|s,a)e^{-\beta V^{\pi}(s^{\prime})}$, we have that
$\mathbb{E}\bigg[\sum_{s^{\prime}}1_{\{X_{n}=s^{\prime}\}}(s^{\prime}|s,a)e^{-\beta V^{\pi}(s^{\prime})}\bigg]$ $=\sum_{s^{\prime}}\mathbb{E}\big[1_{\{X_{n}=s^{\prime}\}}\big]e^{-\beta V^{\pi}(s^{\prime})}$ (63)
$=\sum_{s^{\prime}}P(s^{\prime}|s,a)e^{-\beta V^{\pi}(s^{\prime})}$ (64)
and that it is bounded in $e^{-\beta\frac{1}{1-\gamma}},1]$. Also, since
$\sum_{s^{\prime}}\widehat{P}(s^{\prime}|s,a)e^{-\beta V^{\pi}(s^{\prime})}=\frac{1}{N}\sum_{n=1}^{N}\sum_{s^{\prime}}1_{\{X_{n}=s^{\prime}\}}(s^{\prime}|s,a)e^{-\beta V^{\pi}(s^{\prime})},$ (65)
it follows directly from Hoeffding’s inequality that
$\mathbb{P}\bigg(\bigg|\sum_{s^{\prime}}[P(s^{\prime}|s,a)-\widehat{P}(s^{\prime}|s,a)]e^{-\beta V^{\pi}(s^{\prime})}\bigg|\geq\varepsilon\bigg)\leq 2\exp\bigg(-\frac{2N\varepsilon^{2}}{\big(1-e^{-\beta\frac{1}{1-\gamma}}\big)^{2}}\bigg).$ (66)
Thus, by picking $N=\frac{(1-e^{-\beta\frac{1}{1-\gamma}})^{2}}{2\varepsilon^{2}}\log(2SA/\delta)$ and a union bound,
$\mathbb{P}\bigg(\max_{s,a}|\sum_{s^{\prime}}[P(s^{\prime}|s,a)-\widehat{P}(s^{\prime}|s,a)]e^{-\beta V^{\pi}(s^{\prime})}|\geq\varepsilon\bigg)\leq\delta\,.$ (67)
Appendix F Analysis of MB-RS-QVI: Missing Proofs
F.1 Proof of Lemma 1
Proof.
$V(s):=\max_{a}Q(s,a)\,,\qquad\ X(s):=\max_{a}W(s,a)\,.$
Assuming that $\beta>0$ (the case $\beta<0$ is completely similar), we then have
$\|Q^{\prime}-W^{\prime}\|$ $=Q^{\prime}(s,a)-W^{\prime}(s,a)$
$=\gamma\|V-X\|$
$\leq\gamma\|Q-W\|\,$,
and the lemma follows. ∎
F.2 Proof of Lemma 2
Proof.
Solving $\frac{\gamma^{k}}{1-\gamma}$ for $k$, we get that if $k>\frac{\log(\frac{1}{(1-\gamma)\varepsilon})}{\log(1/\gamma)}$, then $\|Q_{k}-Q^{*}\|<\varepsilon$, thus proving the first claim.
To show the other claim, we start by noting that $\|V^{\pi_{k}}-V^{*}\|\leq\|Q^{\pi_{k}}-Q^{*}\|$. Note also that by design we have that $\mathcal{T}^{\pi_{k}}Q^{\pi_{k}}=Q^{\pi_{k}}$ and that $\mathcal{T}Q_{k}=\mathcal{T}^{\pi_{k}}Q_{k}$. Thus,
$\|Q^{\pi_{k}}-Q^{*}\|\leq\|Q^{\pi_{k}}-Q_{k}\|+\|Q_{k}-Q^{*}\|\,.$ (68)
The first term in the right-hand side is bounded as follows:
$\|Q^{\pi_{k}}-Q_{k}\|$ $=\|\mathcal{T}^{\pi_{k}}Q^{\pi_{k}}-Q_{k}\|$
$\leq\gamma\|Q^{\pi_{k}}-Q_{k}\|+\gamma\|Q_{k}-Q_{k-1}\|\,$,
which means that
$\|Q^{\pi_{k}}-Q_{k}\|\leq\frac{\gamma}{1-\gamma}\|Q_{k}-Q_{k-1}\|\leq\frac{\gamma^{k}}{1-\gamma}\|Q_{1}-Q_{0}\|\leq\frac{\gamma^{k}}{(1-\gamma)^{2}}\,.$ (69)
The proof is completed by observing that picking $k>\frac{\log(\frac{2}{(1-\gamma)^{2}\varepsilon})}{\log(1/\gamma)}$ implies $\|V^{\pi_{k}}-V^{*}\|<\varepsilon$. ∎
F.3 Proof of Lemma 3
Proof.
There are four different cases to consider, namely the combinations arrising from $\beta>0$ vs $\beta<0$ and wether on the state-action pair $(s,a)$ that realizes the maximum it holds that $Q_{1}(s,a)>Q_{2}(s,a)$ or $Q_{2}(s,a)>Q_{1}(s,a)$.
Case 1: $\beta>0,Q_{1}(s,a)>Q_{2}(s,a)$.
$\|Q_{1}-Q_{2}\|$ $=\frac{\gamma}{\beta}\ln\bigg(\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{2}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$=\frac{\gamma}{\beta}\ln\bigg(\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})+\beta(V_{2}(s^{\prime})-V_{1}(s^{\prime}))}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\frac{\gamma}{\beta}\ln\bigg(e^{\beta\|V_{1}-V_{2}\|}\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$=\gamma\|V_{1}-V_{2}\|+\frac{\gamma}{\beta}\ln\bigg(\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\ln\bigg(1+\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}-\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\frac{|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}|}{e^{-\beta\frac{1}{1-\gamma}}}\,.$
Rearranging the terms yields the asserted result:
$\|Q_{1}-Q_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{\beta\frac{1}{1-\gamma}}}{\beta}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}\big|\,.$
Case 2: $\beta>0$ and $Q_{1}(s,a)<Q_{2}(s,a)$. The proof is very similar but the extension $V_{2}(s)=V_{1}(s)+V_{2}(s)-V_{1}(s)$ is now done in the numerator instead:
$\|Q_{1}-Q_{2}\|$ $=\frac{\gamma}{\beta}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{2}(s^{\prime})}}\bigg)$
$=\frac{\gamma}{\beta}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})-\beta(V_{2}(s^{\prime})-V_{1}(s^{\prime}))}}\bigg)$
$\leq\frac{\gamma}{\beta}\ln\bigg(e^{\beta\|V_{1}-V_{2}\|}\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$=\gamma\|V_{1}-V_{2}\|+\frac{\gamma}{\beta}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\ln\bigg(1+\frac{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}-\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{-\beta V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\frac{|\sum_{s^{\prime}}[P_{1}(s^{\prime}|s,a)-P_{2}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}|}{e^{-\beta\frac{1}{1-\gamma}}},$
which again yields
$\|Q_{1}-Q_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{\beta\frac{1}{1-\gamma}}}{\beta}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{-\beta V_{1}(s^{\prime})}\big|\,.$
Case 3: $\beta<0$ and $Q_{1}(s,a)>Q_{2}(s,a)$.
$\|Q_{1}-Q_{2}\|$ $=\frac{\gamma}{|\beta|}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{2}(s^{\prime})}}\bigg)$
$=\frac{\gamma}{|\beta|}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})-|\beta|(V_{2}(s^{\prime})-V_{1}(s^{\prime}))}}\bigg)$
$\leq\frac{\gamma}{|\beta|}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})-|\beta|\|V_{1}-V_{2}\|}}\bigg)$
$=\gamma\|V_{1}-V_{2}\|+\frac{\gamma}{|\beta|}\ln\bigg(\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}\bigg)$
$\leq\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{\beta}\ln\bigg(1+\frac{\sum_{s^{\prime}}P_{1}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}-\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}{\sum_{s^{\prime}}P_{2}(s^{\prime}|s,a)e^{|\beta|V_{1}(s^{\prime})}}\bigg)$
$\gamma\|Q_{1}-Q_{2}\|+\frac{\gamma}{|\beta|}e^{|\beta|\frac{1}{1-\gamma}}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{|\beta|\big[V_{1}(s^{\prime})-\frac{1}{1-\gamma}\big]}\big|,$
which implies
$\|Q_{1}-Q_{2}\|\leq\frac{\gamma}{1-\gamma}\frac{e^{|\beta|\frac{1}{1-\gamma}}}{\beta}\big|\sum_{s^{\prime}}[P_{2}(s^{\prime}|s,a)-P_{1}(s^{\prime}|s,a)]e^{|\beta|[V_{1}(s^{\prime})-\frac{1}{1-\gamma}]}\big|.$
Case 4: $\beta<0$ and $Q_{2}(s,a)>Q_{1}(s,a)$. The proof of this case is similar to the other three cases and is omitted and the final part of the lemma follows by the triangle inequality and the fact that $e^{-|\beta|V_{1}(s)}\leq 1$ and $e^{-|\beta|(\frac{1}{1-\gamma}-V_{1}(s))}\leq 1$. ∎
Appendix G Lower Bound on Bernoulli Likelihood Ratio
In this section, we revisit and develop a technical result that bounds the likelihood ratio of two samples under different hypotheses on a high probability event. Parts of the proof closely resembles parts of Lemma 17 in [23]; however, we stress that our treatment fixes an error in the proof, which however requires slightly stronger assumptions than those imposed in [23]. In addition, while the result in [23] only considers $p\geq\frac{1}{2}$, ours deal with both cases of $p\geq\frac{1}{2}$ and $p<\frac{1}{2}$.
@@ Appendix G Lower Bound on Bernoulli Likelihood Ratio
For $c_{1}=32$, it holds that $\frac{L_{1}(W)}{L_{0}(W)}1_{\mathcal{E}}\geq\frac{2\theta}{c_{2}}1_{\mathcal{E}}$ .
@@
$\log(\frac{c_{2}}{2\theta})=\log\bigg(\frac{c_{2}}{2}\exp\bigg[\frac{c_{1}\alpha^{2}t}{p(1-p)}\bigg]\bigg)\leq\log\bigg(\exp\bigg[\frac{c_{1}\alpha^{2}t}{p(1-p)}\bigg]\bigg)=\frac{c_{1}\alpha^{2}t}{p(1-p)}\,$,
@@
$\frac{L_{1}(W)}{L_{2}(W)}1_{\mathcal{E}}\geq\bigg(\frac{2\theta}{c_{2}}\bigg)^{\frac{2\sqrt{2}}{\sqrt{c_{1}}}+\frac{2(1-p)}{p\cdot c_{1}}+\frac{2}{c_{1}}}1_{\mathcal{E}},$
@@
$\frac{L_{1}(W)}{L_{2}(W)}1_{\mathcal{E}}\geq\frac{2\theta}{c_{2}}1_{\mathcal{E}}\,.$
@@
$\frac{L_{1}(W)}{L_{0}(W)}1_{\mathcal{E}}\geq 1_{\mathcal{E}}\bigg(\frac{2\theta}{c_{2}}\bigg)^{\frac{4}{c_{1}}+\frac{2\sqrt{2}}{\sqrt{c_{1}}}}\geq 1_{\mathcal{E}}\bigg(\frac{2\theta}{c_{2}}\bigg)\,.$
Appendix H Proofs of Lower Bounds
H.1 Lower Bound for Q-value Learning
For a lower bound we construct the following class of MDPs with $S^{\prime}:=S+2$ states and $A$ actions where the first states are labelled $S_{1},,...,s_{S},s^{G},s^{B}$ and the actions are labelled $a_{1},...,a_{A}$. The states $s^{G}$ and $s^{B}$ are absorbing under any actions and $R(s^{G},a)=1$ for all $j$ and $R(s^{B},a)=0$ for all $a\in A$. For the states $s\in\{s_{1},...,s_{S}\}$, we have that $R(s,a)=0$ for all $a\in A$. We have $SA$ state-action pair combinations from $\{s_{1},...,s_{S}\}\times A=:Z$ on which we assume some ordering allowing us to write $z_{i},i\in[SA]$. Finally for all state-action pairs $z_{i}\in[SA]$ we have $P(s^{G}|z_{i})=q_{i}$ and $P(s^{B}|z_{i})=1-q_{i}$ for some $q_{i}\in[0,1]$. The structure of this class of MDPs allows us to get lower bounds on the samples needed to learn the $Q$-value of each state-action pair $z_{i}$ and then use the fact that samples used to learn the $Q$-values for different state-action pairs bring no information on eachother to get the final bound.

### v3

@@ D.3 Proof of Lemma 3
Appendix F Proofs of Lower Bounds
F.1 Lower Bound on Bernoulli Likelihood Ratio
We revisit and develop a technical result that bounds the likelihood ratio of two samples under different hypotheses on a high probability event. Parts of the proof closely resembles parts of Lemma 17 in [30]; however, we stress that our treatment fixes an error in the proof, which however requires slightly stronger assumptions than those imposed in [30]. In addition, while the result in [30] only considers $p\geq\frac{1}{2}$, ours deal with both cases of $p\geq\frac{1}{2}$ and $p<\frac{1}{2}$.
@@ F.1 Lower Bound on Bernoulli Likelihood Ratio
For $c_{1}=32$, it holds that $\frac{L_{1}(W)}{L_{0}(W)}\mathbbmss{1}_{\mathcal{E}}\geq\frac{2\theta}{c_{2}}\mathbbmss{1}_{\mathcal{E}}$ .
@@
$\log\Big(\frac{c_{2}}{2\theta}\Big)=\log\bigg(\frac{c_{2}}{2}\exp\bigg[\frac{c_{1}\alpha^{2}t}{p(1-p)}\bigg]\bigg)\leq\log\bigg(\exp\bigg[\frac{c_{1}\alpha^{2}t}{p(1-p)}\bigg]\bigg)=\frac{c_{1}\alpha^{2}t}{p(1-p)}\,$,
@@
$\frac{L_{1}(W)}{L_{2}(W)}\mathbbmss{1}_{\mathcal{E}}\geq\bigg(\frac{2\theta}{c_{2}}\bigg)^{\frac{2\sqrt{2}}{\sqrt{c_{1}}}+\frac{2(1-p)}{pc_{1}}+\frac{2}{c_{1}}}\mathbbmss{1}_{\mathcal{E}},$
@@
$\frac{L_{1}(W)}{L_{2}(W)}\mathbbmss{1}_{\mathcal{E}}\geq\frac{2\theta}{c_{2}}\mathbbmss{1}_{\mathcal{E}}\,.$
@@
$\frac{L_{1}(W)}{L_{0}(W)}\mathbbmss{1}_{\mathcal{E}}\geq\mathbbmss{1}_{\mathcal{E}}\bigg(\frac{2\theta}{c_{2}}\bigg)^{\frac{4}{c_{1}}+\frac{2\sqrt{2}}{\sqrt{c_{1}}}}\geq\mathbbmss{1}_{\mathcal{E}}\bigg(\frac{2\theta}{c_{2}}\bigg)\,.$
F.2 Lower Bound for Q-value Learning
For a lower bound we construct the following class of MDPs with $S^{\prime}:=S+2$ states and $A$ actions where the first states are labelled $S_{1},,...,s_{S},s^{G},s^{B}$ and the actions are labelled $a_{1},...,a_{A}$. The states $s^{G}$ and $s^{B}$ are absorbing under any actions and $R(s^{G},a)=1$ for all $j$ and $R(s^{B},a)=0$ for all $a\in\mathcal{A}$. For the states $s\in\{s_{1},...,s_{S}\}$, we have that $R(s,a)=0$ for all $a\in\mathcal{A}$. We have $SA$ state-action pair combinations from $\{s_{1},...,s_{S}\}\times\mathcal{A}=:Z$ on which we assume some ordering allowing us to write $z_{i},i\in[SA]$. Finally for all state-action pairs $z_{i}\in[SA]$ we have $P(s^{G}|z_{i})=q_{i}$ and $P(s^{B}|z_{i})=1-q_{i}$ for some $q_{i}\in[0,1]$. The structure of this class of MDPs allows us to get lower bounds on the samples needed to learn the $Q$-value of each state-action pair $z_{i}$ and then use the fact that samples used to learn the $Q$-values for different state-action pairs bring no information on each other to get the final bound.